b/2 - a/3 = 7
b/5 + a/6 = 1

Answer by mananth(16946)

b/2 - a/3 = 7
3b-2a =6*7
3b-2a=42.............1

b/5 + a/6 = 1
6b+5a= 30*1
6b+5a= 30............2
-2a+3b= 42 .............1
5a+6b=30 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 2
8a-12b= -168
10a+12b=60
Add the two equations
18a = -108
/18
a=-6
plug value of a in (1)
-2a+3b= 42
12+3b=42
3b=42-12
3b=30
b=10